OpenStudy (anonymous):

please help me expand the equation using taylor formula f(x)= tan x up to term x^3

OpenStudy (anonymous):

Do you know how to apply Taylor's Theorem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[f(x)=f(a)+(x-a)f'(a)+\frac{(x-a)^2}{2}f''(a)+\frac{(x-a)^3}{6}f'''(a)\\ \] use a=pi/4 for simplicity and find the three derivatives, subtitute for 'a' and expand

OpenStudy (anonymous):

\[f(x)=\tan(x)\\ f'(x)=\sec^2(x)\\ f''(x)=2\tan(x)\sec^2(x) \] and so on